Nebo is a town with a present population of 100,000. The population increases every year by 5%. Determine

what the population size

a) was three years ago

b) will be in two years’ time

c) will be in 7 years’ time

[1] a) 85 000 b) 110 000 c) 135 000

[2] a) 90 000 b) 115 763 c) 140 711

[3] a) 86 384 b) 110 250 c) 140 711

[4] a) 90 703 b) 115 000 c) 135 000

Question 20

After months of negotiations, Jabu’s annual salary was increased by 6.35% to R330 000. Determine his

previous monthly salary and the monthly increase he will be getting.

[1] Original salary of R20 955.00, with an increase of R6 545.00 per month.

[2] Original salary of R25 858.02, with an increase of R1 641.98 per month

[3] Original salary of R27 500.00, with an increase of R1 746.25 per month.

[4] Original salary of R25 858.02, with an increase of R1 746.25 per month


1
Expert's answer
2021-05-13T18:08:12-0400

1. 100,000/1.05^3 = 86,384.

So, the correct answer is [3].

20. The previous salary is 330,000/1.0635/12 = 25,858.02.

The increase in salary is:

25,858.02×0.0635 = 1,641.98.

So, the correct answer is [2].